OpenStudy (skyerose):

Which of the following correctly uses absolute value to show the distance between −70 and 15?

OpenStudy (nevillesnetwork):

What are the provided answers?

jhonyy9 (jhonyy9):

from -70 till zero is 70 plus from zero till 15 is 15 so 70+15 = ?

OpenStudy (nevillesnetwork):

Is this an answer? |−70 − 15| = |−85| = 85 units

OpenStudy (nevillesnetwork):

Because if it is an answer, than it would be that one...

jhonyy9 (jhonyy9):

why - 15 bc. this 15 is in distance plus ?

OpenStudy (nevillesnetwork):

To help us better, you might want to give us the selection of answers...

OpenStudy (nevillesnetwork):

But, from what I can see... the answer would be |−70 − 15| = |−85| = 85 units

jhonyy9 (jhonyy9):

do you mean absolute value of -70 result plus 70 and assumed to 15 = 85 right ?

OpenStudy (rcp031):

Absolute value is defined as distance from zero.